# LLMEval-3: A Large-Scale Longitudinal Study on Robust and Fair Evaluation of Large Language Models

### v3

@@ 3.3 Research Question II
Finding 5: Dynamic rankings diverge significantly from static benchmarks, revealing a different landscape of model capabilities.
To quantify the difference between dynamic and static evaluation, we compare the model rankings from LLMEval-3 with those from two prominent static benchmarks, AGIEval and C-Eval. As shown in Table 3, all correlations are moderate ($\rho\approx 0.66\text{--}0.71$) and statistically insignificant ($p>0.1$), confirming that static evaluations yield markedly different model rankings compared to our dynamic, contamination-resistant framework.
@@
Table 3: Spearman’s rank correlation between LLMEval-3 and static benchmarks.
Finding 6: Static Benchmarks Suffer from Severe Data Contamination.
@@
We observe that on public static benchmarks all models achieve substantially higher completion counts (up to 248), indicating extensive leakage of these questions into the training corpora. In contrast, on our private dataset the number of successful completions drops markedly (maximum 179), with most models near or below 100, demonstrating minimal memorization. This confirms that public static benchmarks suffer from serious data leakage risks, whereas our private question bank—being excluded from large-scale pretraining data—effectively mitigates such risks.
@@
Table 4: Comparison of successful fill-in completions for different models on static benchmarks and LLMEval-3.
@@ 3.4 Research Question III
Finding 7: The relative ranking system demonstrates exceptional stability, with negligible variance across multi-round resampling and varying sample sizes.
To validate the stability of the LLMEval-3 ranking system, we conduct rigorous stress tests involving multi-round resampling and varying sample sizes. In our resampling test, we draw two independent question samples and find that the model ranking order remains identical across both runs, with relative scoring exhibiting remarkably low variance (e.g., $\sigma^{2}=1.68$ for DeepSeek-V3 and just 0.25 for o3-mini). Furthermore, we test the system’s sensitivity to question volume by running evaluations with sample sizes of n=1000, 2000, and 4000. Across all volumes, the ranking order holds firm. These results, detailed in Appendix D, provide strong empirical evidence that our relative ranking system is exceptionally robust and reliable.
Figure 4: Cohen’s $\kappa$ coefficients measuring agreement between human evaluators and three LLM judges across evaluations. GPT-4o achieves almost perfect agreement with human judgments ($\kappa=0.907$).
Finding 8: The relative ranking system via LLM-as-Judge achieves high human-machine agreement and demonstrates superior robustness over alternative ranking systems.
To validate our relative ranking system, we assess it on two critical dimensions: human alignment and methodological robustness. First, we measure its agreement with human experts by conducting two rounds of human evaluation on 13 models and calculating Cohen’s $\kappa$ coefficient against our LLM-as-a-Judge. As shown in Figure 4, the system achieve near-perfect agreement with human judgments, confirming its reliability. Second, we conduct an ablation study comparing our relative ranking against a traditional absolute scoring system (Elo). The results in Figure 5 show that our relative ranking achieves higher Spearman’s $\rho$ correlation and greater stability across varying sample sizes. This dual validation provides strong evidence that our relative ranking system is both aligned with human evaluation and more robust than traditional ranking approaches.
@@ 4 Related Work
Static Knowledge-Intensive Benchmarks such as MMLU (mmlu) and C-Eval (c-eval), provided foundational insights by assessing factual knowledge, reasoning, and cross-domain generalization in language models. These benchmarks established important baselines for measuring model performance across diverse academic domains. However, their static question sets significantly amplify the risk of data leakage, causing inflated evaluation scores and misleading perceptions of model capabilities (DBLP:journals/corr/abs-2412-03597; leakage:xu2024). Moreover, repeated use of fixed datasets induces benchmark-specific overfitting, with models memorizing test answers rather than achieving genuine generalization (leakage:deng2024). LLMEval-1 and LLMEval-2 (llmeval:zhang2024) also fall into this static category. These fundamental vulnerabilities in static evaluation have motivated researchers to explore more dynamic, robust evaluation approaches that better reflect true model generalization abilities.
Dynamic Human-Preference Evaluation such as LMSYS (lmsys:zheng2024), Chatbot Arena (chatbotarena:chiang2024), and AlpacaEval (AlpacaEval) introduce dynamic evaluation settings by leveraging human preferences to compare model outputs in pairwise or multi-turn interactions. These approaches effectively capture conversational quality, coherence, and real-time performance, offering valuable insights beyond traditional static benchmarks. However, they often lack depth in evaluating domain-specific or complex reasoning tasks, and their reliance on non-expert crowdworkers can introduce bias and reduce reliability in assessments of specialized knowledge (humanprefer:raju2024). Furthermore, the focus on conversational abilities may not adequately assess systematic knowledge coverage across academic disciplines.
LLM-as-a-Judge has emerged as a promising paradigm for scalable evaluation, where language models act as evaluators of other models’ outputs. Notable works include G-Eval, which uses GPT-4 with chain-of-thought reasoning and form-filling paradigms to assess natural language generation quality (geval:liu2023), and MT-Bench-101, which evaluates multi-turn dialogues through pairwise comparisons with GPT-4 justifications (mtbench101). These approaches offer significant advantages including high scalability, rapid deployment, cost-effectiveness, and demonstrated alignment with human preferences. However, most existing efforts focus primarily on eliciting better scoring mechanisms, while systematic validation and calibration of the judges, along with mitigation of inherent biases, remain underexplored (judging:zheng2023). Our work addresses these gaps by providing comprehensive analysis and systematic validation of the LLM-as-a-judge paradigm with rigorous human-machine agreement studies.
@@ 5 Conclusion
We introduced LLMEval-3, a dynamic evaluation framework built on a private 220k-question bank, contamination-resistant pipeline, two-layer anti-cheating architecture, and LLM-as-Judge relative ranking to deliver fair, robust, and reproducible LLM assessments. Over a 20-month longitudinal study of nearly 50 open-source and proprietary models, we observed a consistent performance ceiling near 90%, persistent domain-specific gaps in literature, medicine, and military knowledge, and widespread data leakage in static benchmarks. Our relative ranking method showed negligible variance under multi-round resampling and varying sample sizes, while achieving near-perfect agreement with human experts. We further demonstrated that prompting format has minimal impact, whereas external web-search augmentation yields significant gains in knowledge-intensive tasks. These findings confirm the superiority of dynamic, contamination-resistant evaluation over static paradigms and highlight the need for more trustworthy benchmarking practices to guide future LLM development.
@@ Appendix A Dataset
This section provides supplementary information on our LLMEval-3 dataset employed in the study, to clarify the academic disciplines and question types covered, and to elaborate methodologies for scaling the questions.
@@ A.1 Categories of Academic Displines
Categories of Academic Disciplines
@@
Categories of Question Types
The raw dataset encompasses a diverse range of original question formats, including multiple-choices, fill-in-the-blank, true-or-false, short-answer, term explanation, and material analysis questions. Following question expansion and formatting, we have unified all questions and their answers into a fill-in-the-blank-like question-answer format, abandoning the complex answer structures of various question types, such as options A to E in multiple-choice questions, and “true” or “false” in true-or-false questions. This natural question-answer format enables the dataset to more thoroughly showcase the model’s capabilities.
@@ A.2 Details of Expanding the Dataset
Original Data Construction Pipeline
@@
Data Expanding Pipeline
@@
Original Multiple-Choices questions: Title: It is known that the sixth level of a complete binary tree (let the root be the first level) has eight leaves, then the number of nodes of the complete binary tree is at most ( ). A. 39 B. 52 C. 111 D. 119 Answer: C Expanded questions: Question 1: It is known that the sixth level of a complete binary tree (let the root be the first level) has eight leaves, then the number of nodes of the complete binary tree is at most: ( ) Is it correct to place the answer “39” in the provided space? Answer: False Question 2: It is known that the sixth level of a complete binary tree (let the root be the first level) has eight leaves, then the number of nodes of the complete binary tree is at most: ( ) Is it correct to place the answer “52” in the provided space? Answer: False Question 3: It is known that the sixth level of a complete binary tree (let the root be the first level) has eight leaves, then the number of nodes of the complete binary tree is at most: ( ) Is it correct to place the answer “111” in the provided space? Answer: True Question 4: It is known that the sixth level of a complete binary tree (let the root be the first level) has eight leaves, then the number of nodes of the complete binary tree is at most: ( ) Is it correct to place the answer “119” in the provided space? Answer: False
@@ Appendix B Prompts
This section presents the complete set of prompts used in LLMEval-3 for different evaluation paradigms. We provide the specific prompt templates for few-shot learning, chain-of-thought reasoning, and LLM-based automated evaluation to ensure reproducibility of our experimental results.
@@
Input: Here are several examples: Question: Please determine the correctness of the following statement. Answer with true/false and provide a reason. If all elements below the diagonal in the adjacency matrix of a directed graph are zero, then the graph must have a topological ordering. Answer: true. Because if all elements below the diagonal in the adjacency matrix of a directed graph are zero, the graph is a Directed Acyclic Graph (DAG), so it must have a topological ordering. Question: Please explain the following term: “Double Hundred Policy”. Answer: It refers to “let a hundred flowers bloom, let a hundred schools of thought contend.” This was a policy on science and culture officially proposed by Mao Zedong in 1956 and confirmed by the Central Committee of the Communist Party of China. The policy was severely undermined after 1957, but was reestablished and implemented following the Third Plenary Session of the 11th Central Committee. Question: Please determine the correctness of the following statement. Answer with true/false and provide a reason. According to the convertibility theory, the scope of commercial banks’ assets expanded from short-term turnover loans to consumer loans. Answer: false. Convertibility Theory: also known as the asset conversion theory. This theory suggests that to maintain liquidity for withdrawals, commercial banks can invest part of their funds in transferable securities. Since these profitable assets can be sold at any time and converted into cash, loans are not necessarily limited to short-term and self-liquidating types. Clearly, this theory emerged in the context of developing financial instruments and markets. Significance: it expanded the scope of bank asset operations. Drawbacks: it does not guarantee that assets can be liquidated without capital loss (which requires high asset quality and stable market conditions); it is also constrained by central bank monetary policy (e.g., the risk of a rise in discount rates). This theory provides a theoretical basis for Chinese commercial banks to engage in securities business (investment operations). However, in China, commercial banks’ investment activities are restricted due to: 1. limited investment instruments and underdeveloped credit mechanisms; 2. management systems narrowing investment scopes, and separation of operations preventing commercial banks from investing in stocks; 3. the nature of state-owned commercial banks limits their willingness for autonomous investment. Question: Why can the results of animal experiments not be fully applied to clinical practice? Answer: Because there are differences between humans and animals not only in cellular morphology and metabolism, but also fundamentally due to the highly developed human nervous system, which is associated with language and thought (the second signaling system). Although there are similarities, the essential differences mean that human diseases cannot all be replicated in animals. Even if they can be replicated, animal responses are simpler than human responses. Therefore, results from animal experiments cannot be mechanically and fully applied to clinical practice without analysis. Only by comparing, analyzing, and synthesizing animal experiment results with clinical data can they be used as references in clinical medicine and provide a basis for studying the causes, mechanisms, prevention, and treatment of clinical diseases. The following is the question to be answered: Question:{question} Answer:
@@
Input: Please evaluate the following response from the LLM regarding a discipline-specific question based on the following criteria. You must score it on a scale of 0, 1, 2 or 3 stars: Overall Rating: 0 star indicates wrong answer with a wrong explanation 1 stars indicate wrong answer but a partially reasonable explanation 2 stars indicate a correct answer with a partially reasonable explanation 3 stars indicate an correct answer with a reasonable explanation User: {question} LLM: {LLM response} The correct answer to user’s question is: {correct answer} You must provide your feedback in the following format: “Overall Rating”:numbers of its stars(int) The reason why you gave this rating: ¡Your Reason¿(str)’
@@
Appendix C LLMEval-3 Leaderboard

### v5

@@ 3.3 Research Question II
3.3.0.1 Finding 5: Dynamic rankings differ from static benchmarks, revealing a distinct view of model capabilities.
To quantify the discrepancy between dynamic and static evaluation, we compare the model rankings from LLMEval-Fair with two representative static benchmarks, AGIEval (AGIEval) and C-Eval (c-eval). As shown in Table 3, the rank correlations are moderate ($\rho\approx 0.65\text{--}0.72$), indicating that static benchmark rankings do not consistently align with those produced by our dynamic evaluation.
@@
Table 3: Spearman’s rank correlation between LLMEval-Fair and static benchmarks.
3.3.0.2 Finding 6: Static Benchmarks Suffer from Severe Data Contamination.
@@
We observe that on public static benchmarks almost all models achieve substantially higher completion counts, indicating extensive leakage of these questions into the training corpora. In contrast, on our private dataset, completion counts drop markedly, with most models near or below 100. Overall, these results suggest that public static benchmarks may suffer from significant leakage risk, whereas our private question bank, which is not included in large-scale pretraining corpora, supports a more contamination-resistant evaluation.
@@
Table 4: Comparison of successful fill-in completions for different models on static benchmarks and LLMEval-Fair.
@@ 3.4 Research Question III
3.4.0.1 Finding 7: The relative ranking system demonstrates exceptional stability, with negligible variance across multi-round resampling and varying sample sizes.
To validate the stability of the LLMEval-Fair ranking system, we conduct rigorous tests involving multi-round resampling and varying sample sizes. In our resampling test, we draw two independent question samples and find that the model ranking order remains identical across both runs, with relative scoring exhibiting remarkably low variance (e.g., $\sigma^{2}=1.68$ for DeepSeek-V3 and just 0.25 for o3-mini). Furthermore, we test the system’s sensitivity to question volume by running evaluations with sample sizes of n=1000, 2000, and 4000. Across all volumes, the ranking order holds firm. These results, detailed in Appendix F, provide strong empirical evidence that our relative ranking system is exceptionally robust and reliable.
3.4.0.2 Finding 8: The relative ranking system via LLM-as-Judge achieves high human-machine agreement and demonstrates superior robustness over alternative ranking systems.
To validate our relative ranking system, we assess it on two critical dimensions: human alignment and methodological robustness. First, we measure its agreement with human experts by conducting two rounds of human evaluation and calculating Cohen’s $\kappa$ coefficient against our LLM-as-a-Judge. As shown in Figure 5, the system achieve near-perfect agreement with human judgments, confirming its reliability. Second, we conduct an ablation study comparing our relative ranking against an Elo-style rating baseline implemented under identical evaluation conditions. Specifically, both methods are evaluated on the same fixed pool of 1,000 queries, and the curves in Figure 5 are obtained by repeatedly sub-sampling from this pool to vary the effective sample size. The results show that our relative ranking consistently achieves higher Spearman’s $\rho$ correlation and exhibits greater stability as the sample size changes. This dual validation provides strong evidence that our relative ranking system is both aligned with human evaluation and more robust than conventional rating-based alternatives.
Figure 4: Cohen’s $\kappa$ coefficients measuring agreement between human evaluators and three LLM judges across evaluations. GPT-4o achieves almost perfect agreement ($\kappa=0.907$).
@@ 4 Related Work
Static Knowledge-Intensive Benchmarks like MMLU (mmlu) and C-Eval (c-eval) measure factual knowledge and reasoning, yet their fixed question sets risk data leakage—inflating scores and misrepresenting capabilities (DBLP:journals/corr/abs-2412-03597; leakage:xu2024)—while promoting overfitting where models memorize test answers instead of generalizing (leakage:deng2024). This static paradigm, shared by LLMEval-1 and LLMEval-2 (llmeval:zhang2024), is now driving research toward more dynamic and robust evaluation methods.
Dynamic Human-Preference Evaluation (e.g., LMSYS (lmsys:zheng2024), Chatbot Arena (chatbotarena:chiang2024), AlpacaEval (AlpacaEval)) uses human preferences from pairwise or multi-turn interactions to assess conversational quality and real-time performance beyond static benchmarks. However, these methods often lack depth in domain-specific reasoning, rely on non-expert annotators prone to bias in specialized assessments (humanprefer:raju2024), and focus on dialogue at the expense of systematic academic knowledge coverage.
LLM-as-a-Judge employs language models as scalable evaluators of model outputs, as seen in G-Eval (geval:liu2023) and MT-Bench-101 (mtbench101). This paradigm offers high scalability, low cost, and strong human preference alignment. However, existing work focuses heavily on scoring mechanisms, leaving systematic judge calibration and bias mitigation underexplored (judging:zheng2023). Our work addresses this by rigorously analyzing and validating the paradigm through human–machine agreement studies.
@@ 5 Conclusion
We introduced LLMEval-Fair, a dynamic, contamination-resistant evaluation framework built on a private 220k-question bank, a two-layer anti-cheating architecture, and an LLM-as-Judge relative ranking pipeline. A 30-month longitudinal study of nearly 60 open-source and proprietary models revealed a consistent performance ceiling near 90%, systematic gaps in literature, medicine, and military knowledge, and widespread data leakage in static benchmarks. Our relative ranking method demonstrated negligible variance under multi-round resampling with varying sample sizes and achieved near-perfect agreement with human experts. We further confirmed that prompting format has minimal impact on performance in knowledge-intensive tasks, underscoring the superiority of dynamic, contamination-resistant evaluation over static benchmarks and the need for more trustworthy benchmarking practices.
Limitations
The main limitation is the sheer size of the question bank, about 220,000 items, which makes a truly comprehensive evaluation resource-intensive. Running large-scale inference over the full set, validating and aggregating results, and maintaining the dataset through cleaning, deduplication, and difficulty calibration require substantial compute, time, and human effort. As a result, this evaluation is more feasible for teams with stable infrastructure and ongoing operational capacity, while smaller groups may struggle to reproduce an equally complete assessment with the same rigor and cadence.
Appendix
@@ Appendix A Dataset
This section provides supplementary information on our LLMEval-Fair dataset employed in the study, to clarify the academic disciplines and question types covered, and to elaborate methodologies for scaling the questions.
@@ A.1 Categories of Academic Displines
A.1.1 Categories of Academic Disciplines
@@
A.1.2 Categories of Question Types
The raw dataset encompasses a diverse range of original question formats, including multiple-choices, fill-in-the-blank, true-or-false, short-answer, term explanation, and material analysis questions. Following question expansion and formatting, we have unified all questions and their answers into a fill-in-the-blank-like question-answer format, abandoning the complex answer structures of various question types, such as options A to E in multiple-choice questions, and “true" or “false" in true-or-false questions. This natural question-answer format enables the dataset to more thoroughly showcase the model’s capabilities.
@@ A.2 Details of Expanding the Dataset
A.2.1 Original Data Construction Pipeline
@@
A.2.2 Data Expanding Pipeline
@@
Original Multiple-Choices questions: Title: It is known that the sixth level of a complete binary tree (let the root be the first level) has eight leaves, then the number of nodes of the complete binary tree is at most ( ). A. 39 B. 52 C. 111 D. 119 Answer: C Expanded questions: Question 1: It is known that the sixth level of a complete binary tree (let the root be the first level) has eight leaves, then the number of nodes of the complete binary tree is at most: ( ) Is it correct to place the answer “39" in the provided space? Answer: False Question 2: It is known that the sixth level of a complete binary tree (let the root be the first level) has eight leaves, then the number of nodes of the complete binary tree is at most: ( ) Is it correct to place the answer “52" in the provided space? Answer: False Question 3: It is known that the sixth level of a complete binary tree (let the root be the first level) has eight leaves, then the number of nodes of the complete binary tree is at most: ( ) Is it correct to place the answer “111" in the provided space? Answer: True Question 4: It is known that the sixth level of a complete binary tree (let the root be the first level) has eight leaves, then the number of nodes of the complete binary tree is at most: ( ) Is it correct to place the answer “119" in the provided space? Answer: False
@@ Appendix B Prompts
This section presents the complete set of prompts used in LLMEval-Fair for different evaluation paradigms. We provide the specific prompt templates for few-shot learning, chain-of-thought reasoning, and LLM-based automated evaluation to ensure reproducibility of our experimental results.
@@
Input:
Here are several examples:
Question:
Please determine the correctness of the following statement. Answer with true/false and provide a reason. If all elements below the diagonal in the adjacency matrix of a directed graph are zero, then the graph must have a topological ordering.
Answer:
true. Because if all elements below the diagonal in the adjacency matrix of a directed graph are zero, the graph is a Directed Acyclic Graph (DAG), so it must have a topological ordering.
Question:
Please explain the following term: “Double Hundred Policy".
Answer:
It refers to “let a hundred flowers bloom, let a hundred schools of thought contend." This was a policy on science and culture officially proposed by Mao Zedong in 1956 and confirmed by the Central Committee of the Communist Party of China. The policy was severely undermined after 1957, but was reestablished and implemented following the Third Plenary Session of the 11th Central Committee.
Question:
Please determine the correctness of the following statement. Answer with true/false and provide a reason. According to the convertibility theory, the scope of commercial banks’ assets expanded from short-term turnover loans to consumer loans.
Answer:
@@
false. Convertibility Theory: also known as the asset conversion theory. This theory suggests that to maintain liquidity for withdrawals, commercial banks can invest part of their funds in transferable securities. Since these profitable assets can be sold at any time and converted into cash, loans are not necessarily limited to short-term and self-liquidating types. Clearly, this theory emerged in the context of developing financial instruments and markets. Significance: it expanded the scope of bank asset operations. Drawbacks: it does not guarantee that assets can be liquidated without capital loss (which requires high asset quality and stable market conditions); it is also constrained by central bank monetary policy (e.g., the risk of a rise in discount rates). This theory provides a theoretical basis for Chinese commercial banks to engage in securities business (investment operations). However, in China, commercial banks’ investment activities are restricted due to: 1. limited investment instruments and underdeveloped credit mechanisms; 2. management systems narrowing investment scopes, and separation of operations preventing commercial banks from investing in stocks; 3. the nature of state-owned commercial banks limits their willingness for autonomous investment.
Question:
Why can the results of animal experiments not be fully applied to clinical practice?
Answer:
Because there are differences between humans and animals not only in cellular morphology and metabolism, but also due to the highly developed human nervous system, which is associated with language and thought (the second signaling system). Although there are similarities, the essential differences mean that human diseases cannot all be replicated in animals. Even if they can be replicated, animal responses are simpler than human responses. Therefore, results from animal experiments cannot be mechanically and fully applied to clinical practice without analysis. Only by comparing, analyzing, and synthesizing animal experiment results with clinical data can they be used as references in clinical medicine and provide a basis for studying the causes, mechanisms, prevention, and treatment of clinical diseases.
The following is the question to be answered:
Question:{question}
Answer:
@@
Input: Please evaluate the following response from the LLM regarding a discipline-specific question based on the following criteria. You must score it on a scale of 0, 1, 2 or 3 stars: Overall Rating: 0 star indicates wrong answer with a wrong explanation 1 stars indicate wrong answer but a partially reasonable explanation 2 stars indicate a correct answer with a partially reasonable explanation 3 stars indicate an correct answer with a reasonable explanation User: {question} LLM: {LLM response} The correct answer to user’s question is: {correct answer} You must provide your feedback in the following format: “Overall Rating":numbers of its stars(int) The reason why you gave this rating: <Your Reason>(str)’
@@
Appendix C Evaluation Model Selection
To ensure a comprehensive and rigorous assessment, we conducted a preliminary evaluation on a broad set of 59 large language models. From this extensive pool, we selected a representative subset of 17 models for the detailed analysis presented in the main text. The detailed release dates for these selected models are provided in Table 7.
The proprietary frontier models include the OpenAI lineup, featuring the widely deployed GPT-4o and GPT-4o-search (openai2024gpt4o), alongside the next-generation flagship GPT-5 (openai2025gpt5). Furthermore, we assess the o1 series (openai2024o1) and the efficient o3-mini (openai2025o3mini), which represent specialized reasoning models trained with large-scale reinforcement learning to solve complex scientific and mathematical problems.
We include models capable of extended reasoning from other major providers. Anthropic is represented by the standard Claude-Sonnet-4 (anthropic2024claude4), as well as the Claude-Sonnet-4-Thinking and the advanced 4.5-Thinking variants (anthropic2025claude45), which operate in a specialized mode to perform self-reflection before generating responses. Similarly, Google’s contribution consists of Gemini-2.5-Pro (google2025gemini25) and its reasoning-enhanced counterpart, Gemini-2.5-Pro-Thinking (google2025gemini25), which incorporates internal chain-of-thought processes.
Regarding open-weights architectures and diverse scaling strategies, the study includes the DeepSeek family: DeepSeek-V3 (deepseek2025v3), a strong Mixture-of-Experts (MoE) model, and DeepSeek-R1 (deepseek2025r1), a model specifically optimized for reasoning tasks through post-training. The Qwen-3 series is also evaluated to represent distinct points on the parameter spectrum, featuring both the massive 235B model and the compact 32B variant (alibaba2025qwen3).
Finally, we evaluate other high-performing systems to ensure a broad representation of the current landscape. This includes Moonshot’s Kimi-K2 (moonshot2025kimik2) and the Doubao-1.5 series. Notably, Doubao-1.5-Thinking-Pro (bytedance2025doubao) serves as a key reference in our study, having demonstrated state-of-the-art capabilities in preliminary screenings.
Model Released Date Model Released Date
Doubao-1.5-Thinking-Pro (bytedance2025doubao) Apr 15, 2025 Claude-Sonnet-4.5-Thinking (anthropic2025claude45) Sep 29, 2025
Doubao-1.5-Pro (bytedance2025doubao) Jan 15, 2025 Claude-Sonnet-4-Thinking (anthropic2024claude4) May 14, 2025
Gemini-2.5-Pro-Thinking (google2025gemini25) Jun 5, 2025 Claude-Sonnet-4 (anthropic2024claude4) May 14, 2025
Gemini-2.5-Pro (google2025gemini25) Jun 5, 2025 o1 (openai2024o1) Dec 17, 2024
DeepSeek-R1 (deepseek2025r1) May 28, 2025 o3-mini (openai2025o3mini) Jan 29, 2025
DeepSeek-V3 (deepseek2025v3) Mar 24, 2024 GPT-4o-search (openai2024gpt4o) Nov 20, 2024
Qwen-3-235B (alibaba2025qwen3) Apr 29, 2025 GPT-4o (openai2024gpt4o) Nov 20, 2024
Qwen-3-32B (alibaba2025qwen3) Apr 29, 2025 Kimi-K2 (moonshot2025kimik2) Sep 5, 2025
GPT-5 (openai2025gpt5) Jun 7, 2025
Table 7: The representative list of models selected from a total of 59 evaluated LLMs. The list is arranged in two columns to optimize space. Dates for unreleased models are estimated based on technical previews.
Appendix D The Elo Rating System
The Elo rating system is a widely recognized method for calculating relative skill levels in zero-sum games, originally developed for chess and recently popularized for evaluating Large Language Models (e.g., Chatbot Arena). This system derives ratings from the outcomes of pairwise comparisons, providing a probabilistic framework to predict the likelihood of one entity outperforming another. Given two entities $A$ and $B$ with current ratings $R_{A}$ and $R_{B}$, the expected score $E_{A}$ (representing the probability of $A$ winning) is calculated using a logistic curve:
$E_{A}=\frac{1}{1+10^{(R_{B}-R_{A})/400}}$ (3)
Following a match, the ratings are updated based on the discrepancy between the actual outcome $S_{A}$ (where 1 represents a win, 0 a loss, and 0.5 a tie) and the expected probability. The update rule is given by $R_{A}^{\prime}=R_{A}+K(S_{A}-E_{A})$, where $K$ is a constant factor that determines the sensitivity of the rating adjustment.
Appendix E LLMEval-Fair Leaderboard